# On finiteness properties of separating semigroup of real curve

### v2

@@ 2. Deleting points from separating divisors
Lemma 1 (Lemma 3.2 in [^1] In [9] the ($\Leftarrow$) part of Lemma 3.2. assumes the divisor is non-special. In fact, this assumption can be removed.[9]).
Let $p_{1},\ldots,p_{n}\in\mathbb{R}X$ be pairwise distinct points. Then, the divisor $P=p_{1}+\ldots+p_{n}$ is separating if and only if there exist a tuple of positive (with respect to some complex orientation) tangent vectors $v_{i}\in T_{p_{i}}(\mathbb{R}X),\ i=1,\ldots,n$ such that
@@
Note that Condition ($\dagger$1) can be reformulated as follows. The tuple $v=(v_{1},\ldots,v_{n})$ defines a tangent vector to $\operatorname{Sym}^{n}(X)$ at the point $P$. Let $\varphi\colon\operatorname{Sym}^{n}(X)\to\operatorname{Pic}^{n}(X)$ denote the Abel-Jacobi map. Then Condition ($\dagger$1) is precisely the statement that $v\in\operatorname{Ker}(d_{P}\varphi)$.
@@ 4. Proof of the main theorem
As the known examples show (see, e.g., item (3) in § 3.3), the bound $\deg P_{i}\leq 4g-3$ is far from sharp.
@@ Acknowledgements
This paper was inspired by a question posed to me by Sergei O. Gorchinskiy after my talk at the Vth Conference of Mathematical Centers of Russia, concerning the description of the separating semigroup in finite terms. Further development of this work, and, in particular, the proof of the main theorem was stimulated by a question from Olivier Benoist, which I received after the publication of the first version of the preprint. I am grateful to both of them for these insightful questions.

### v3

@@ 2. Deleting points from separating divisors
Lemma 1 (Lemma 3.2 in [^1] In [9] the ($\Leftarrow$) part of Lemma 3.2 assumes that the divisor is non-special. In fact, this assumption is redundant.[9]).
Let $p_{1},\ldots,p_{n}\in\mathbb{R}X$ be pairwise distinct points. Then, the divisor $P=p_{1}+\ldots+p_{n}$ is separating if and only if there exists a tuple of positive (with respect to some complex orientation) tangent vectors $v_{i}\in T_{p_{i}}(\mathbb{R}X),\ i=1,\ldots,n$ such that
@@
Proof.
First note that Condition ($\dagger$1) can be reformulated as follows. The tuple $v=(v_{1},\ldots,v_{n})$ defines a tangent vector to $\operatorname{Sym}^{n}(X)$ at $P$. Let $\varphi\colon\operatorname{Sym}^{n}(X)\to\operatorname{Pic}^{n}(X)$ denote the Abel-Jacobi map. Then Condition ($\dagger$1) is precisely the statement that $v\in\operatorname{Ker}(d_{P}\varphi)$. So, the implication $(\Rightarrow)$ follows directly from the Abel-Jacobi theorem.
Let us prove the implication $(\Leftarrow)$. Again, it follows from the Abel-Jacobi theorem that there exists a smooth deformation
$p_{i}(t)\colon[0,1]\to\mathbb{R}X\colon\quad P(t)=\sum_{i=1}^{n}p_{i}(t)\in|P|,\quad p_{i}(0)=p_{i},\quad\frac{d}{dt}\bigg|_{t=0}\ p_{i}(t)=v_{i}.$
Since all $v_{i}$ are positive with respect to some complex orientation on $\mathbb{R}X$, for sufficiently small $t\in[0,1]$ all points $p_{i}(t)$ move in the direction of the complex orientation. Hence for $t_{0}\ll 1$ divisors $Q=P(t_{0})$ and $P$ interlace (that is, every component of $\mathbb{R}X\setminus Q$ contains exactly one point of $P$ and vice versa). Since $Q\in|P|$, there exists a meromorphic function $f$ such that (f)=PQ(f)=P-Q. Since $Q$ and $P$ interlace, by [7, Lemma 2.10] the meromorphic function $f$ is a separating morphism. ∎
@@ 4. Proof of the main theorem
As the known examples show (see, e.g., item (3) in § 3.3), the bound $\deg P_{i}\leq 4g-3$ is far from sharp. This leads us to the following question.
@@ Acknowledgements
This paper was inspired by a question posed to me by Sergei O. Gorchinskiy after my talk at the Vth Conference of Mathematical Centers of Russia, concerning the description of the separating semigroup in finite terms. Further development of this work, and, in particular, the proof of the main theorem was stimulated by a question from Olivier Benoist, which I received after the publication of the first version of the preprint. I am grateful to both of them for these questions.